Circular No. 8300

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ap IN PGC 29306 Further to IAUC 8299, J. Graham and W. Li report the LOSS dis-

covery, on KAIT images taken on Mar. 8.31 and 9.24 UT, of an apparent supernova (mag ∼ 16.5) located at α = 10^{h}05^{m}43^{s}.81, δ = +10^{o}16′17′′.1 (equinox 2000.0), which is 10′′.5 west and 19′′.5 south of the nucleus of PGC 29306. A KAIT image taken on Feb. 10.41 showed nothing at this position (limiting mag ∼ 19.0).

PSR 1259–63 M. Beilicke, Universität Hamburg; M. Ouchrif, Laboratoire de Phy-

sique Nucléaire de Haute Energie, Université Paris VI and VII; G. Rowell, Max-Planck-Institut für Kernphysik; and S. Schlenker, Humboldt-Univer-sität, Berlin; on behalf of the ‘High Energy Stereoscopic System Collaboration’, report detection of very-high-energy γ-ray emission from the binary millisecond pulsar PSR 1259–63 (1236.72-day orbit) above a threshold of ∼ 200 GeV (in the observed zenith-angle range, 40^{o}–45^{o}) at the 5.9σ level, where previously upper limits were reported by the CANGAROO collaboration (Kawachi et al. 2004, Ap.J., in press). The source was observed for a total of 4.6 hr live-time between Feb. 26 and Mar. 5 with the full four-telescope Cherenkov array. Preliminary estimates yield a flux at ∼ 10 percent of the level of the Crab nebula (0.4 photons/min γ-ray excess detected after selection cuts). Periastron passage of the pulsar was expected on Mar. 7.43 UT. Observations at other wavelengths are strongly encouraged, especially on Mar. 20–25, when the neutron star will cross the inclined Be disk for the second time. Particle acceleration associated with the pulsar wind is predicted to result in observable GeV/TeV emission also after periastron (Kirk et al. 1999, Astroparticle Phys. 10, 31).

COMETS C/2003 T5, C/2003 T6, C/2003 T7 (SOHO) Further to IAUC 8291, K. McGleam reports her measures of three

additional Kreutz sungrazing comets found on SOHO C3 website images by J. Sachs. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-E28.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 T5 Oct. 8.929  12^{h}33.^{m}6  − 7^{o}21′
C/2003 T6  9.471  12 42.3  − 7 37
C/2003 T7  9.471  12 28.7  − 7 58
2004 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green